OpenStudy (anonymous):

2 quick problems (: Solve for r: C = 2^3.14r Solve the equation 2 thirds (x − 2) = 4x for x

OpenStudy (anonymous):

ln C = 3.14r ln 2 r = ln C / 3.14 ln 2 (2/3)x - 4/3 = 4x 10x / 3 = -4/3 can you complete these?